OpenStudy (anonymous):

how many solutions does 5x + 3x − 4 = 10 have

OpenStudy (freckles):

do you know how to solve it? Also If you solve a linear equation ax+b where a isn't 0 equals a constant function c; such as ax+b=c when you solve for x we take the following steps ax=c-b subtracted b on both sides x=(c-b)/a divided a on both sides how many solutions do you see?

OpenStudy (anonymous):

im not sure i get it.